OpenStudy (anonymous):

find the sum of 1/8 + 1/6 =i think the denominator is 48

OpenStudy (sasogeek):

denominator is 24.

OpenStudy (anonymous):

how do i find the sum

OpenStudy (sasogeek):

3+4/24=7/24

OpenStudy (anonymous):

\[1/8 + 1/6=3/24+4/24=(3+4)/24=7/24\]

OpenStudy (anonymous):

1/8 =1/6 = 0.291